# MNSC-0001-2024.00

### v1

@@
Surety bonds are financial agreements between a contractor (principal) and obligee (project owner) to complete a project. However, most large-scale projects involve multiple contractors, creating a network and introducing the possibility of incomplete obligations to propagate and result in project failures. Typical models for risk assessment assume independent failure probabilities within each contractor. However, we take a network approach, modeling the contractor network as a directed graph where nodes represent contractors and project owners and edges represent contractual obligations with associated financial records. To understand risk propagation throughout the contractor network, we extend the celebrated Friedkin-Johnsen model and introduce a stochastic process to simulate principal failures across the network. From a theoretical perspective, we show that under natural monotonicity conditions on the contractor network, incorporating network effects leads to increases in both the average risk and the tail probability mass of the loss distribution (i.e. larger right-tail risk) for the surety organization. We further use data from a partnering insurance company to validate our findings, estimating an approximately 2% higher exposure when accounting for network effects.
@@ 1 Introduction
Surety bonds are a foundational mechanism in contractual risk management, widely used to guarantee the completion of projects in sectors such as construction, infrastructure, and public works. In a typical surety agreement, the surety company guarantees to the obligee (the project owner) that the principal (a contractor) will fulfill the terms of a bonded contract. If the principal fails to perform, the surety must step in to ensure project completion, often absorbing substantial financial losses in the process (russell1990surety). These agreements are not only mandated for public contracts under laws such as the U.S. Miller Act of 1935 (uscode40_3131), but also play a growing role in private-sector project financing (wambach2011surety). Despite their ubiquity, surety bonds remain difficult for contractors to secure, in part due to the unexpected and systemic nature of failures, which can leave insurers liable; for instance, there were over \$21 billion in claims between 1990 and 1997 in the U.S. alone (wambach2011surety).
@@
The pricing of surety bonds depends critically on the ability to assess the default risk of individual contractors. A large body of work in finance and insurance focuses on this task, estimating failure probabilities using firm-level covariates such as credit ratings, leverage ratios, and liquidity metrics (kim2019default). While such models (often using statistical machine learning) have improved the accuracy of idiosyncratic risk estimation, they share a key limitation: they assume failures occur independently across firms. This assumption neglects an increasingly salient feature of real-world contracting environments: network dependencies among contractors, subcontractors, and project owners (see Fig. 1 for a toy contractor network, later in Section 5 we consider a real-world contractor network with $\sim 30,000$ organizations). In practice, most large-scale projects involve multi-tiered contractual relationships, where the performance of one firm is contingent on the timely execution of work by others. As highlighted by recent industry reviews (assuredpartners2024surety, bci2021supplychain), subcontractor failure is one of the leading causes of bonded losses. For instance, if a plumbing contractor cannot begin work until the electrical subcontractor completes their portion of a build, the default of the latter creates a domino effect. In such environments, risk is not merely a function of a firm’s own characteristics, but also of its position in the broader contractor network. Moreover, these cascading effects are pervasive across domains. In supply chains, for example, the bankruptcy of an upstream supplier can cripple downstream production. In collaborative research, a delay in one lab’s work can stall the entire study. In interbank lending networks, financial contagion spreads through credit exposures. Across all these settings, a network-aware perspective is essential to understand and mitigate systemic risk.
@@
However, despite growing recognition of these interdependencies, there remains a lack of formal models that account for network effects in surety risk assessment. While prior work has examined dynamic models of credit contagion and equilibrium default (benzoni2015modeling, nickerson2017debt), these frameworks are often not tailored to the structure of bonded contractor networks, where obligations are directional. Notably, recent theoretical work in stochastic dynamics and mean-field models (amini2022dynamic, carmona2013mean) explores related ideas, but typically assumes irreversible failures or continuous-time evolution, and does not address the specific distributional shape of systemic losses, a critical consideration for insurers concerned about tail risk. Our work is motivated by this important gap. Namely, we seek to answer the following research questions:
@@ 1.1 Main Contributions
In an attempt to tackle these questions, one of the main contributions of this work is a network-based modeling framework for risk propagation in surety-based contractor networks. Traditional models like kim2019default assume independent failure probabilities, not capturing how failures spread through contractor relationships. Instead, we represent the contracting environment as a directed network $G=(\mathcal{V},\mathcal{E})$, where nodes correspond to principals (contractors) and obligees (project owners), and edges capture the potential flow of risk through a contract for bonded work from principal to obligee. In contrast to network-unaware models of failure dynamics, we introduce a stochastic process $\mathbf{X}^{t}=(X_{i}^{t})_{i\in\mathcal{V}}$ which represents whether node $i\in\mathcal{V}$ in the network fails at timestep $t\in\mathbb{N}$. This stochastic process evolves according to the following simple dynamics (see Eq. 1):
$X_{i}^{t+1}=\textsf{Bernoulli}\Big((1-\alpha_{i})r_{i}+\alpha_{i}\sum_{j\in\delta_{\text{in}}(i)}w_{ij}X_{j}^{t}\Big),$
@@
To understand how risk propagates in the network, we analyze the stationary distribution of the stochastic failure process $X_{i}^{t}$. We start off by showing that the marginal failure probabilities $m_{i}^{t}=\mathbb{E}\mathopen{}\mathclose{{\left[X_{i}^{t}}}\right]$ converge to a unique fixed point $m_{i}$ as their limiting failure probability (Section 3.1). This result generalizes the Friedkin-Johnsen model of opinion dynamics to a setting with heterogeneous bias parameters. We further quantify the rate of convergence, establishing an exponential decay in $|m_{i}^{t}-m_{i}|$ governed by the operator norm of a squared weight-adjusted adjacency matrix, and additionally show that the convergence occurs in finite time for acyclic contractor networks (Fig. 3).
Building on the mean-field analysis, we next study the full joint distribution over failures. The stochastic process $X_{i}^{t}$ defines a Markov chain over $\{0,1\}^{n}$, where $n$ is the number of nodes (contractors and obligees) in the network. However, naive analysis will establish its convergence to the stationary distribution in $\mathcal{O}(2^{n})$ time (resnick2013adventures). In contrast, we show that in acyclic graphs, the convergence occurs in at most $d$ steps, where $d>0$ is the length of the longest path in the network (Definition 4.4). For general (not necessarily acyclic) networks, we leverage the structure of the stochastic process to show it admits a coupling. This allows us to develop a contraction-based analysis, and show that the rate of convergence scales logarithmically with respect to $n$ (Section 4.1). Together, these results characterize the stationary behavior of the failure stochastic process, and show that the stationary distribution can be simulated efficiently.
@@
Our modeling framework allows us to quantify how systemic risk is amplified by the network structure beyond what traditional independent risk models predict. The main insight of our analysis is we show that when obligees hire riskier contractors on average (Section 3.3), which is often observed in contractor networks (dietz2018mitigating), both the expected failure probabilities (Section 3.3) and the right-tail probability mass of the loss distribution (Section 4.2) increase over time. To formalize this, we design a monotone coupling between the highly-correlated stochastic process to its mean-field, allowing us to show that these risk measures are stochastically dominated across time. We further outline conditions (Remark 3.2) such that the total risk and loss in the network is strictly larger (Remark 3.2). With this analysis we observe that contractor failures are not uniformly impactful—intermediary nodes, which serve as bridges between multiple principals and obligees, are the primary drivers behind the propagation of project incompletion across the network. In essence, by accounting for network effects, we see that surety organizations may be fundamentally underestimating how risky a contractor network is. Lastly, we define an eigenvalue centrality (Definition 3.1) that captures the extent to which a contractor’s risk influences the broader network through downstream intermediaries. To summarize, our results formally characterize when and how network structure exacerbates risk.
@@ 1.2 Related Literature
Our work lies at the intersection of operations research, economics, applied probability, and network analysis, with close connections to models of financial contagion and surety risk. See caccioli2018network for a broad survey.
@@
Surety bonds are a widely adopted mechanism to protect against contractor defaults and offer several advantages over traditional insurance (schubert2002point). Surety bonds are typically priced using firm-level information such as financial ratios or credit data (schubert2002point, kim2019default). A large empirical literature applies statistical and machine learning methods, including logit regression (tserng2014prediction), SVMs (tserng2011svm, horta2013company), ensemble learning (choi2018predicting), and Bayesian networks (cao_2022), to estimate individual contractors’ default probabilities from accounting data (barboza2017machine, nguyen2025bankruptcy, shumway2001forecasting, vassalou2004default). Although some models incorporate macroeconomic covariates (shumway2001forecasting, vassalou2004default), all of this literature treats contractors as independent units, with no mechanism by which one contractor’s failure propagates to others.
In practice, however, systemic factors and subcontracting dependencies create correlations in defaults. Historical events such as the 1980s oil embargo led to widespread contractor failures despite strong individual credit profiles (russell1990surety), and modern construction projects often hinge on “lower-tier” subcontractors, whose disruptions can cascade through the supply chain (dietz2018mitigating, bci2021supplychain). Motivated by these limitations, our model augments contractor-level default estimates from the existing literature with network interactions, capturing how contractual ties generate correlated risks and amplify potential losses.
@@
Much of the contagion literature studies interbank lending. An early contribution is allenFinancialContagion2000, who model contagion as an equilibrium phenomenon in interbank markets, where small liquidity shocks can spread through overlapping claims. Subsequent work often uses threshold models in which a node defaults once losses from its neighbors exceed a threshold (watts2002simple, gai2010contagion, elliott2014financial, acemoglu2015systemic). These frameworks highlight how localized shocks can spread systemically, but their dynamics are typically deterministic and tied to an initial shock event. More recent analyses consider noisy threshold contagion, where a small probability of below-threshold adoption can accelerate the spread of complex contagion (ecklesLongTiesAccelerate2024). Unlike threshold contagion models that assume diminishing returns from additional affected neighbors, our model is stochastic and cumulative: a contractor’s failure risk grows as an aggregate function of weighted neighbor defaults and an idiosyncratic baseline.
Other extensions consider multilayer contagion with mutations, where new strains can emerge as the process spreads and heterogeneity across layers (e.g., schools, workplaces) shapes transmission (SoodSridhar2023). This underscores how ignoring heterogeneity in either the contagion type or the network structure can miscalculate systemic risk. In our setting, the analogous challenge is to capture heterogeneity in contractor obligations and recovery, rather than multilayer or mutating contagions.
A further distinction is that most threshold models assume that failed institutions remain insolvent, whereas in surety settings, defaults trigger intervention from the surety organization to ensure project completion. Models with recovery or stochasticity include recovery in reinsurance networks (klages2020cascading), Gaussian noise in asset values (ramirez2023stochastic), and dynamic link formation with Cramér–Lundberg premiums/claims (amini2022dynamic). However, these approaches rely on explicit thresholds or detailed balance-sheet information uncommon for bonded contractors. Instead, we propose a stochastic cascade model with heterogeneous exposures and explicit recovery, tailored to contractor–surety networks.
davis2001infectious introduce an alternative to threshold contagion by modeling “infectious” defaults, where bonds fail independently or through Bernoulli contagion within a sector. This framework shares similarities with our approach in that defaults can arise either idiosyncratically or via neighbors, but it assumes uniform exposures and fully connected networks. Moreover, it lacks recovery, which are central in the surety context. Our model builds on this probabilistic contagion idea while incorporating heterogeneity in network structure, firm characteristics, and explicit recovery.
@@
Opinion dynamics provides a natural basis for modeling failure cascades, where contractors correspond to individuals, and default probabilities correspond to opinions. A central feature is neighbor influence, allowing local shocks to generate global effects across the network. For instance, benzoni2015modeling show how investors’ beliefs about bond pricing can propagate through financial networks in ways that resemble contagion. A well-known framework is the Friedkin–Johnsen model, in which agents reconcile their intrinsic beliefs with their neighbors’ views (FriedkinJohnsen). Our model has a similar structure: defaults may be triggered by neighbors but are also shaped by inherent contractor-level failure probabilities. In fact, the mean-field of our stochastic process corresponds to a variation of the Friedkin–Johnsen model when opinions are reinterpreted as default probabilities (see Section 3.1). Related extensions, such as the interacting Pólya Urn model of tang2024estimating, further highlight how noisy observations and social pressure interact with intrinsic beliefs.
@@
The Friedkin–Johnsen model is closely connected to eigenvector-based centrality measures such as Bonacich centrality and PageRank. Namely, bonacich2001eigenvector highlight that Bonacich centrality can be applied to the same situations as the Friedkin-Johnsen model, with the distinction that the latter is concerned with the limiting state, while Bonacich centrality quantifies the level of influence each node has on the final equilibrium. This centrality is analogous to the equilibrium output in response to a shock in the input-output model introduced by leontief1986input and appears in other settings in the economics literature. For example, acemoglu2012network studies a network of production sectors and shows that the volatility of aggregate output scales with the size of an eigenvector-based centrality vector that is closely related to Bonacich centrality. A special class of the Friedkin-Johnsen model also coincides with the teleportation model of random surfing introduced by page1999pagerank, proskurnikov2016pagerank.
In financial settings, battiston2012debtrank propose DebtRank, an eigenvector-based measure of systemic importance in interbank lending. While their measure prevents risk from being transmitted multiple times along cycles, in surety networks repeated impacts of earlier failures are precisely what matter. For this reason, we use an eigenvector-based centrality as a complement to our stochastic model, capturing how individual risks amplify through contractual ties (see Section 3.2).

### v2

@@
Surety bonds are financial agreements between a contractor (principal) and obligee (project owner) to complete a project. However, most large-scale projects involve multiple contractors, creating a network and introducing the possibility of incomplete obligations to propagate and result in project failures. Typical models for risk assessment assume independent failure probabilities within each contractor. However, we take a network approach, modeling the contractor network as a directed graph where nodes represent contractors and project owners and edges represent contractual obligations with associated financial records. To understand risk propagation throughout the contractor network, we extend the celebrated Friedkin-Johnsen model and introduce a stochastic process to simulate principal failures across the network. From a theoretical perspective, we show that under natural monotonicity conditions on the contractor network, incorporating network effects leads to increases in the average risk for the surety organization. We further use data from a partnering insurance company to validate our findings, estimating an approximately 2% higher exposure when accounting for network effects.
@@ 1 Introduction
Surety bonds are a foundational mechanism in contractual risk management, widely used to guarantee the completion of projects in sectors such as construction, infrastructure, and public works. In a typical surety agreement, the surety company guarantees to the obligee (the project owner) that the principal (a contractor) will fulfill the terms of a bonded contract. If the principal fails to perform, the surety must step in to ensure project completion, often absorbing substantial financial losses in the process (Russell 1990). These agreements are not only mandated for public contracts under laws such as the U.S. Miller Act of 1935 (United States Code 2002), but also play a growing role in private-sector project financing (Wambach and Engel 2011). Despite their ubiquity, surety bonds remain difficult for contractors to secure, in part due to the unexpected and systemic nature of failures, which can leave insurers liable; for instance, there were over \$21 billion in claims between 1990 and 1997 in the U.S. alone (Wambach and Engel 2011).
@@
The pricing of surety bonds depends critically on the ability to assess the default risk of individual contractors. A large body of work in finance and insurance focuses on this task, estimating failure probabilities using firm-level covariates such as credit ratings, leverage ratios, and liquidity metrics (Kim et al. 2019). While such models (often using statistical machine learning) have improved the accuracy of idiosyncratic risk estimation, they share a key limitation: they assume failures occur independently across firms. This assumption neglects an increasingly salient feature of real-world contracting environments, the presence of network dependencies among contractors, subcontractors, and project owners (see Fig. 1 for a toy contractor network, later in Section 5 we consider a real-world contractor network with $\sim 30,000$ organizations). In practice, most large-scale projects involve multi-tiered contractual relationships, where the performance of one firm is contingent on the timely execution of work by others. As highlighted by recent industry reviews (Brouse 2024, Business Continuity Institute 2021), subcontractor failure is one of the leading causes of bonded losses. For instance, if a plumbing contractor cannot begin work until the electrical subcontractor completes their portion of a build, the default of the latter creates a domino effect. In such environments, risk is not merely a function of a firm’s own characteristics, but also of its position in the broader contractor network. Moreover, these cascading effects are pervasive across domains. In supply chains, for example, the bankruptcy of an upstream supplier can cripple downstream production. In collaborative research, a delay in one lab’s work can stall the entire study. In interbank lending networks, financial contagion spreads through credit exposures. Across all these settings, a network-aware perspective is essential to understand and mitigate systemic risk.
@@
However, despite growing recognition of these interdependencies, there remains a lack of formal models that account for network effects in surety risk assessment. While prior work has examined dynamic models of credit contagion and equilibrium default (Benzoni et al. 2015, Nickerson and Griffin 2017), these frameworks are often not tailored to the structure of bonded contractor networks, where obligations are directional. Notably, recent theoretical work in stochastic dynamics and mean-field models (Amini et al. 2022, Carmona et al. 2013) explores related ideas, but typically assumes irreversible failures or continuous-time evolution. Our work is motivated by this important gap. Namely, we seek to answer the following research questions:
@@ 1.1 Main Contributions
In an attempt to tackle these questions, one of the main contributions of this work is a network-based modeling framework for risk propagation in surety-based contractor networks. Traditional models like Kim et al. (2019) assume independent failure probabilities, not capturing how failures spread through contractor relationships. Instead, we represent the contracting environment as a directed network $G=(\mathcal{V},\mathcal{E})$, where nodes correspond to principals (contractors) and obligees (project owners), and edges capture the potential flow of risk through a contract for bonded work from principal to obligee. In contrast to network-unaware models of failure dynamics, we introduce a stochastic process $\mathbf{X}^{t}=(X_{i}^{t})_{i\in\mathcal{V}}$ which represents whether node $i\in\mathcal{V}$ in the network fails at timestep $t\in\mathbb{N}$. This stochastic process evolves according to the following simple dynamics (see Eq. 1):
$X_{i}^{t+1}=\textsf{Bernoulli}\Big((1-\alpha_{i})r_{i}+\alpha_{i}\sum_{j\in\delta_{\text{in}}(i)}w_{ij}X_{j}^{t}\Big).$
@@
To understand how risk propagates in the network, we analyze the stationary distribution of the stochastic failure process $X_{i}^{t}$. We start off by showing that the marginal failure probabilities $m_{i}^{t}=\mathbb{E}\mathopen{}\mathclose{{\left[X_{i}^{t}}}\right]$ converge to a unique fixed point $m_{i}$ as their limiting failure probability (Section 3.1). This result generalizes the Friedkin-Johnsen model of opinion dynamics to a setting with heterogeneous bias parameters (Friedkin and Johnsen 1990). We further quantify the rate of convergence, establishing an exponential decay in $|m_{i}^{t}-m_{i}|$ governed by the operator norm of a squared weight-adjusted adjacency matrix, and additionally show that the convergence occurs in finite time for acyclic contractor networks (Fig. 3).
Building on the mean-field analysis, we next study the full joint distribution over failures. The stochastic process $X_{i}^{t}$ defines a Markov chain over $\{0,1\}^{n}$, where $n$ is the number of nodes (contractors and obligees) in the network. However, naive analysis will establish its convergence to the stationary distribution in $\mathcal{O}(2^{n})$ time (Resnick 2013). In contrast, we show that in acyclic graphs, the convergence occurs in at most $d$ steps, where $d>0$ is the length of the longest path in the network (Definition 4.4). For general (not necessarily acyclic) networks, we leverage the structure of the stochastic process to show it admits a coupling. This allows us to develop a contraction-based analysis, and show that the rate of convergence scales logarithmically with respect to $n$ (Section 4.1). Together, these results characterize the stationary behavior of the failure stochastic process, and show that the stationary distribution can be simulated efficiently.
@@
Our modeling framework allows us to quantify how systemic risk is amplified by the network structure beyond what traditional independent risk models predict. The main insight of our analysis is we show that when obligees hire riskier contractors on average (Section 3.3), which is often observed in contractor networks (Dietz 2018), the expected failure probabilities (Section 3.3) increase over time. We further outline conditions (Remark 3.2) such that the total risk and loss in the network is strictly larger (Remark 3.2). With this analysis we observe that contractor failures are not uniformly impactful, intermediary nodes, which serve as bridges between multiple principals and obligees, are the primary drivers behind the propagation of project incompletion across the network. In essence, by accounting for network effects, we see that surety organizations may be fundamentally underestimating how risky a contractor network is. Lastly, we define an eigenvalue centrality (Definition 3.1) that captures the extent to which a contractor’s risk influences the broader network through downstream intermediaries. To summarize, our results formally characterize when and how network structure exacerbates risk.
@@ 1.2 Related Literature
Our work lies at the intersection of operations research, economics, applied probability, and network analysis, with close connections to models of financial contagion and surety risk. See Caccioli et al. (2018) for a broad survey.
@@
Surety bonds are a widely adopted mechanism to protect against contractor defaults and offer several advantages over traditional insurance (Schubert and Duke 2002). Surety bonds are typically priced using firm-level information such as financial ratios or credit data (Schubert and Duke 2002, Kim et al. 2019). A large empirical literature applies statistical and machine learning methods, including logit regression (Tserng et al. 2014), SVMs (Tserng et al. 2011, Horta and Camanho 2013), ensemble learning (Choi et al. 2018), and Bayesian networks (Cao et al. 2022), to estimate individual contractors’ default probabilities from accounting data (Barboza et al. 2017, Nguyen et al. 2025, Shumway 2001, Vassalou and Xing 2004). Although some models incorporate macroeconomic covariates (Shumway 2001, Vassalou and Xing 2004), all of this literature treats contractors as independent units, with no mechanism by which one contractor’s failure propagates to others.
In practice, however, systemic factors and subcontracting dependencies create correlations in defaults. Historical events such as the 1980s oil embargo led to widespread contractor failures despite strong individual credit profiles (Russell 1990), and modern construction projects often hinge on “lower-tier” subcontractors, whose disruptions can cascade through the supply chain (Dietz 2018, Business Continuity Institute 2021). Motivated by these limitations, our model augments contractor-level default estimates from the existing literature with network interactions, capturing how contractual ties generate correlated risks and amplify potential losses.
@@
Much of the contagion literature studies interbank lending. An early contribution is Allen and Gale (2000), who model contagion as an equilibrium phenomenon in interbank markets, where small liquidity shocks can spread through overlapping claims. Subsequent work often uses threshold models in which a node defaults once losses from its neighbors exceed a threshold (Watts 2002, Gai and Kapadia 2010, Elliott et al. 2014, Acemoglu et al. 2015). These frameworks highlight how localized shocks can spread systemically, but their dynamics are typically deterministic and tied to an initial shock event. More recent analyses consider noisy threshold contagion, where a small probability of below-threshold adoption can accelerate the spread of complex contagion (Eckles et al. 2024). Unlike threshold contagion models that assume diminishing returns from additional affected neighbors, our model is stochastic and cumulative. A contractor’s failure risk grows as an aggregate function of weighted neighbor defaults and an idiosyncratic baseline.
Other extensions consider multilayer contagion with mutations, where new strains can emerge as the process spreads and heterogeneity across layers (e.g., schools, workplaces) shapes transmission (Sood et al. 2023). This underscores how ignoring heterogeneity in either the contagion type or the network structure can miscalculate systemic risk. In our setting, the analogous challenge is to capture heterogeneity in contractor obligations and recovery, rather than multilayer or mutating contagions.
A further distinction is that most threshold models assume that failed institutions remain insolvent, whereas in surety settings, defaults trigger intervention from the surety organization to ensure project completion. Models with recovery or stochasticity include recovery in reinsurance networks (Klages-Mundt and Minca 2020), Gaussian noise in asset values (Ramirez et al. 2023), and dynamic link formation with Cramér–Lundberg premiums/claims (Amini et al. 2022). However, these approaches rely on explicit thresholds or detailed balance-sheet information uncommon for bonded contractors. Instead, we propose a stochastic cascade model with heterogeneous exposures and explicit recovery, tailored to contractor–surety networks.
Davis and Lo (2001) introduce an alternative to threshold contagion by modeling “infectious” defaults, where bonds fail independently or through Bernoulli contagion within a sector. This framework shares similarities with our approach in that defaults can arise either idiosyncratically or via neighbors, but it assumes uniform exposures and fully connected networks. Moreover, it lacks recovery, which are central in the surety context. Our model builds on this probabilistic contagion idea while incorporating heterogeneity in network structure, firm characteristics, and explicit recovery.
@@
Opinion dynamics provides a natural basis for modeling failure cascades, where contractors correspond to individuals, and default probabilities correspond to opinions. A central feature is neighbor influence, allowing local shocks to generate global effects across the network. For instance, Benzoni et al. (2015) show how investors’ beliefs about bond pricing can propagate through financial networks in ways that resemble contagion. A well-known framework is the Friedkin–Johnsen model, in which agents reconcile their intrinsic beliefs with their neighbors’ views (Friedkin and Johnsen 1990). Our model has a similar structure where defaults may be triggered by neighbors but are also shaped by inherent contractor-level failure probabilities. In fact, the mean-field of our stochastic process corresponds to a variation of the Friedkin–Johnsen model when opinions are reinterpreted as default probabilities (see Section 3.1). Related extensions, such as the interacting Pólya Urn model of Tang et al. (2024), further highlight how noisy observations and social pressure interact with intrinsic beliefs.
@@
The Friedkin–Johnsen model is closely connected to eigenvector-based centrality measures such as Bonacich centrality and PageRank. Namely, Bonacich and Lloyd (2001) highlight that Bonacich centrality can be applied to the same situations as the Friedkin-Johnsen model, with the distinction that the latter is concerned with the limiting state, while Bonacich centrality quantifies the level of influence each node has on the final equilibrium. This centrality is analogous to the equilibrium output in response to a shock in the input-output model introduced by Leontief (1986) and appears in other settings in the economics literature. For example, Acemoglu et al. (2012) studies a network of production sectors and shows that the volatility of aggregate output scales with the size of an eigenvector-based centrality vector that is closely related to Bonacich centrality. A special class of the Friedkin-Johnsen model also coincides with the teleportation model of random surfing introduced by Page et al. (1999), Proskurnikov et al. (2016).
In financial settings, Battiston et al. (2012) propose DebtRank, an eigenvector-based measure of systemic importance in interbank lending. While their measure prevents risk from being transmitted multiple times along cycles, in surety networks repeated impacts of earlier failures are precisely what matter. For this reason, we use an eigenvector-based centrality as a complement to our stochastic model, capturing how individual risks amplify through contractual ties (see Section 3.2).